OpenStudy (maddiie):

Find the solution to each system by the substitution method. 10x+3y=-8 2x+y=2

OpenStudy (anonymous):

second equation gives \[y=2-2x\] substitute in the first equation to get \[10x+3(2-2x)=-8\]

OpenStudy (maddiie):

10x=6x+6=-8

OpenStudy (anonymous):

\[10x+6-6x=-8\] \[4x+6=-8\] \[4x=-8-6=-14\] \[x=-\frac{14}{4}=-\frac{7}{2}\]

OpenStudy (maddiie):

Ok

OpenStudy (anonymous):

now you have to substitute back to find y

OpenStudy (anonymous):

\[y=2-2x\] \[y=2-2\times- \frac{7}{2}\] \[y=2+7\] \[y=9\]

OpenStudy (anonymous):

hold on let me check

OpenStudy (maddiie):

I was doing while you were teaching me and i got the smae thing useless i'm wrong too. I am understanding this

OpenStudy (anonymous):

oh yeah it is right!